The father is 35 the son is 9. How many years will it take the son to be 3 times his fathers age. I have two different answers that I do not know if they are correct. I have tried 35-9= 26X3 = 78. I have also tried 35x3= 105/9 =11.667.
Found 2 solutions by oberobic, stanbon:
Answer by oberobic(2304)

35+x = 3(9+x)
35+x = 27 + 3x
-2x = -8
.
x = 4
.
In 4 years, the father will be 39 and the son will be 13. 39 = 3*13.
.
It is impossible for a son to be 3 times the father's age ever, so I solved it the other way.
The father will be 3 times the son's age in 4 years.
.
Done.

Answer by stanbon(75887)

The father is 35 the son is 9. How many years will it take the son to be 3 times his fathers age.
---------
Note: The son will never be older than the father.
Equation:
Let the number of years be "x".
(9+x) = 3(35+x)
9+x = 115 = 3x
2x = - 106
---
x = -53
----
Father's age: 35-53 = -18
Son's age: 9-53 = -44
Nonsense
=======================
Cheers,
Stan H.
===============